# GFlowGR: Fine-tuning Generative Recommendation Frameworks with Generative Flow Networks

### v1

@@ 4.2. Main Results
Comparison with On-policy RLFT. On-policy strategies (especially reward-based PPO) fail to achieve competitive results for GR. This may stem from out-of-vocabulary item identifiers: LLMs pretrained on general language tasks struggle to explore the OOV item space without a well-defined reward model, leading to poor PPO performance. GRPO alleviates this dependency on the reward model via relative advantages. (Shao et al., 2024)
Comparison with Off-policy RLFT. Vanilla DPO (Rafailov et al., 2023) fails to stably improve upon SFT, particularly on large-scale Yelp data, due to its use of fixed and randomly sampled “Rejected” items. In contrast, SPRec (Gao et al., 2025a) (using self-play sampling) and IPA (Deng et al., 2025) (leveraging collaborative knowledge) enhance DPO through dynamic, guided negative sampling. Among DPO-based methods, S-DPO (Chen et al., [n. d.]), which is capable of incorporating multiple negative samples, achieves the best results and competes closely with GFlowGR-DB.
Figure 3. Parameter study on $K$ and $\lambda$ with TIGER on Beauty. The red dashed line denotes the performance of SFT, which falls outside the plotted range for the figures corresponding to $K$.
@@ 4.4. Ablation Status
In this section, we validate the critical components of GFlowGR: the curriculum-style adaptive trajectory sampler, comprehensive reward model, and uncertainty-based trajectory reweighting. Table 2 presents ablation results for: all reward signal combinations (six variants, first block of the table, where the name of the variant denotes used reward dignals), removal of adaptive sampling (w/o Ada), omission of Augmented Trajectories (w/o Traj), and removal of reweighting (w/o Wt). Key findings include: (a) Loss of any reward signal (augmentation signals, collaborative scores, or token similarities) degrades performance, confirming the necessity of the comprehensive reward model. Notably, $r_{a}$ has the most significant impact, while $r_{c}$ contributes marginally due to CM producing less accurate predictions than GR (Rajput et al., 2023; Wang et al., 2024a), leaving to noisy $r_{c}$ signals. (b) Removing adaptive sampling or using random sampling reduces effectiveness due to unstable training or low-quality augmented samples. (c) Uncertainty-based reweighting contributes to alleviating exposure bias and improving model robustness, validating its role in stabilizing training.
Table 2. Ablation study with TIGER on Beauty.
Metrics $r_{a}$ $r_{c}$ $r_{\text{sim}}$ $r_{a}+r_{c}$ $r_{a}+r_{\text{sim}}$ $r_{c}+r_{\text{sim}}$ w/o Ada w/o Traj w/o Wt GFlowGR
R@5 0.0403 0.0399 0.0404 0.0399 0.0406 0.0381 0.0386 0.0397 0.0388 0.0407
R@10 0.0641 0.0631 0.0634 0.0629 0.0648 0.0600 0.0600 0.0617 0.0615 0.0651
N@5 0.0261 0.0263 0.0260 0.0264 0.0260 0.0253 0.0257 0.0261 0.0254 0.0266
N@10 0.0341 0.0338 0.0343 0.0337 0.0344 0.0323 0.0326 0.0332 0.0327 0.0344
4.5. In-depth Analysis
Case Study. We visualize a case in Figure 4 to study GFlowGR’s potential in addressing exposure bias. The example shows GFlowGR recommending a diverse set of items, including ‘Shampoo’, ‘Scrub’, and ‘Cleanser’ with the ‘Tomatoes’ flavor (bold green), which is a new taste for the user. While SFT focuses on ‘Carrots’-flavored ‘Shampoo’ and ‘Conditioner’ from the same brand, which the user recently purchased (bold orange and red). This suggests GFlowGR alleviates exposure bias by exploring diverse trajectories, generating exploratory results with different flavors and product.
Diversity. For this experiment, we analyze diversity in top-10 recommendations from GFlowGR and SFT across two dimensions. (a) Generation Distribution: We compare models’ item recommendation probabilities against a normal distribution using KL divergence (rather than a uniform distribution because item popularities are inherently different for recommendations (Chen et al., 2023)). GFlowGR achieves a closer match (KL=3.13 vs. SFT’s 3.24). In addition, the generation probability over all items has a 2.2% smaller standard deviation with GFlowGR, indicating a smoother, more diverse distribution. (b) Collaborative Scores: We calculate the collaborative score for the recommended items. The result shows that GFlowGR’s recommended items have 0.4% lower average collaborative scores but a broader range (0.45 vs. SFT’s 0.42) and 3.6% higher standard deviation, showing it prioritizes non-mainstream items while offering greater score diversity. These results demonstrate GFlowGR’s enhanced ability to balance relevance and novelty compared to SFT.
Figure 4. Case study on user-2089 from Beauty with TIGER: We visualized three items from the top-10 list of each method for the brand ‘Yes to’, where the top-1 result of GFlowGR is correct.
@@ 5. Related Work
Generative Recommendations. Existing GR research can be categorized into three directions: pipeline design, item tokenizer innovations, and LLM decoding enhancement. For pipeline design, TIGER (Rajput et al., 2023) pioneered out-of-vocabulary item tokenization and GR workflows, while LC-Rec (Zheng et al., 2024) and ColaRec (Wang et al., 2024b) enhance alignment via tasks or graph collaborative models, with extensions to multi-behavior scenarios (Liu et al., 2024a). The second category focuses on plug-in tokenizers (e.g., LETTER (Wang et al., 2024a) for collaborative embeddings) to boost arbitrary GR frameworks. The third category optimizes decoding (Lin et al., 2025c; Ding et al., 2024). However, LLM fine-tuning for GR remains underexplored, with DPO-based IPA (Deng et al., 2025) as a rare example. The proposed GFlowGR is the first to study the exposure bias and diversity.
@@ 6. Conclusion
In this paper, we propose GFlowGR, a GFlowNets-based fine-tuning framework for GR. Specifically, we design an adaptive trajectory sampler that leverages collaborative knowledge to select augmented trajectories, enabling the model to explore diverse item spaces. We also introduce a comprehensive reward model that integrates multi-faceted signals, including augmentation status, collaborative scores, and token similarities, to provide nuanced feedback for trajectory evaluation. During fine-tuning, GFlowGR extends the vanilla SFT loss by incorporating flow-matching objectives, with trajectory-specific losses reweighted based on uncertainty. By adaptive sampling, heuristic reweighting, and flow-matching targets, GFlowGR enhances recommendation diversity and addresses the exposure bias challenge inherent in GR frameworks. Extensive experimental results validate the effectiveness of our approach across multiple datasets, backbones, and metrics.

### v2

@@ 4.2. Main Results
Comparison with On-policy RLFT. On-policy strategies fail to achieve competitive results for GR. This may stem from their sensitivity to the data. Although GRPO can also learn from item set with relative advantages. (Shao et al., 2024), it maximizes the probability for the positive item, failing to concept the out-of-set items. While GFlowGR assigns the probability proportional to the reward without forcing the sum of the item set as 1, possessing the better robustness in term of the augmented item set.
Comparison with Off-policy RLFT. Vanilla DPO (Rafailov et al., 2023) fails to stably improve upon SFT, particularly on large-scale Yelp data, due to its use of fixed and randomly sampled “Rejected” items. In contrast, SPRec (Gao et al., 2025a) (using self-play sampling) and IPA (Deng et al., 2025) (leveraging collaborative knowledge) enhance DPO through dynamic, guided negative sampling. Among DPO-based methods, S-DPO (Chen et al., [n. d.]), which is capable of incorporating multiple negative samples, achieves the best results and competes closely with GFlowGR-DB. This also highlight the urgent need for the ability of learning from the item set rather than single positive item or item pairs.
@@ 4.4. Ablation Status
In this section, we validate the critical components of GFlowGR: the trajectory sampler and the reward model. Table 3 presents ablation results for: all reward signal combinations (six variants, first block of the table, where the name of the variant denotes used reward dignals), removal of CM-based adaptive sampling (w/o Ada, using random sampling), omission of Augmented Trajectories (w/o Traj), i.e., $N=1$. Key findings include: (a) Loss of any reward signal (augmentation signals, collaborative scores, or token similarities) degrades performance, confirming the necessity of the comprehensive reward model. Notably, $r_{a}$ has the most significant impact, while $r_{c}$ contributes marginally due to CM producing less accurate predictions than GR (Rajput et al., 2023; Wang et al., 2024a), leaving to noisy $r_{c}$ signals. (b) Using random sampling reduces effectiveness due to unstable training or low-quality augmented samples.
4.5. Deployment
Since May 2025, GFlowGR has been fully deployed to train our GR models across all Taobao search advertising platforms, including Taobao Mobile, Xianyu, and Taobao Web (as illustrated in Figure 4).
Offline evaluations, summarized in Table 4, compare GFlowGR against GRPO and SFT on a massive dataset of hundreds of millions of daily requests. To maintain confidentiality, training costs are presented relative to the SFT baseline. While GRPO’s cost is computationally infeasible, GFlowGR requires approximately double the training time of SFT. We note this can be mitigated by scaling resources; doubling the number of GPUs brings GFlowGR’s training time on par with SFT’s. In terms of performance, the metrics mirror our findings on public datasets: GFlowGR achieves a significantly larger improvement over the SFT baseline than GRPO does.
The method’s real-world value was validated in a large-scale online A/B test involving over 100 million user sessions. The test revealed statistically significant improvements in core business metrics, including a 0.43% increase in total revenue. This lift was particularly pronounced in off-homepage search scenarios, which saw a 1.11% revenue increase. These production metrics, collected over 15 consecutive days and verified by Taobao’s internal monitoring systems, confirm the substantial real-world impact.
Figure 4. Deployed business scenarios in Taobao, illustrated with the query “Dress” in the top left red box. Advertisements are indicated by blue boxes on the bottom right of each product. Due to privacy policy, sensitive information has been obfuscated.
Method H@20 N@20 Trianing Cost
[…]
GRPO 0.427 0.408 -
[…]
Table 4. Offline evaluations on production data.
@@ 5. Related Work
Generative Recommendations. According to the composition of GR, existing research can be generally divided into three categories: design of the whole pipeline, innovations on the item tokenizer, and LLM decoding enhancement. The representative works for the first category include TIGER (Rajput et al., 2023), LC-Rec (Zheng et al., 2024), and ColaRec (Wang et al., 2024b). TIGER (Rajput et al., 2023) is the pioneer in exploring expressing items with out-of-vocabulary tokens and constructs the basic GR pipeline, including the training and inference with item tokenizer and LLMs. LC-Rec (Zheng et al., 2024) further enhances the model efficacy by incorporating diverse aligning tasks to align item identifiers to recommendation objectives. Similarly, ColaRec (Wang et al., 2024b) introduces the graph-based collaborative models to align the semantics with recommendations. Some researchers also explore GR frameworks for other recommendation problems, such as multi-behavior recommendations (Liu et al., 2024a). The second category focuses on inventing plug-in item tokenizers to enhance the performance of arbitrary GR frameworks (Zhu et al., 2024; Zheng et al., 2025a; Liu et al., 2024c; Lin et al., 2025b). For example, LETTER (Wang et al., 2024a) introduces the collaborative item embeddings and diversity constraints to control the item tokenization process. While the last category notices the decoding process for GR and designs the decoding for specific recommendation targets (Lin et al., 2025c; Ding et al., 2024). While there are few efforts spent on the LLM fine-tuning specific to GR besides the optimization algorithms for general language-based recommendations (Chen et al., [n. d.]; Gao et al., 2025a). IPA from OneRec (Deng et al., 2025) is the only LLM fine-tuning work specific to GR. The proposed GFlowGR is the pioneer in exploring LLM fine-tuning with the objective of learning from item set.
@@ 6. Conclusion
In this paper, we propose GFlowGR, a GFlowNets-based fine-tuning framework for GR. Specifically, we design a trajectory sampler with four strategies to augment the training trajectories, enabling the model to explore diverse item spaces. We also introduce a reward model that can integrate multi-faceted signals, exemplified with interaction signal, estimated scores, and bussiness target (illustraed as similarities), to provide nuanced feedback for trajectory evaluation. During fine-tuning, GFlowGR extends the vanilla SFT loss by incorporating flow-matching objectives. With the ability of learning from the item set, GFlowGR enhances recommendation diversity and quality. Extensive experimental results validate the effectiveness of our approach across multiple datasets, backbones, and metrics.